Condense the expression to the logarithm of a single quantity.
2 ln 5 + 2 ln z

Found 2 solutions by lynnlo, ikleyn:
Answer by lynnlo(4176)
Answer by ikleyn(53765)

.
Condense the expression to the logarithm of a single quantity.
2 ln 5 + 2 ln z
~~~~~~~~~~~~~~~~~~~~~~~~~~~


2*ln(5) + 2*ln(z) = ln(5^2) + ln(z^2) = ln((5^2)*(z^2)) = ln(25z^2).             ANSWER


The answer in the post by @lynnlo is incorrect.